Prove the above integral.

4185-2039_3899_indefinite6.bmp


Prove the above integral.

Grade:12

1 Answers

Ramesh V
14 years ago

put : x2 + 1/x2 = t

        (2x  - 2)/x3 .dx = dt

         2(x4-1)x3.dx=dt

I = integral of { (x4-1)/x2*x(x2+1+(1/x2)) .dx  }

   = integral of { (1/2.(t)1/2 ).dt }

   = 1/2 * 2*(t)1/2 + C

   = (x2 + (1/x2) +1)1/2  + C

--